OpenStudy (anonymous):

factorise x^2 - y^2 - 2x + 6y - 8 pls give steps for the solution..

OpenStudy (anonymous):

=(x^2-2x+1)-(y^2-6y+9)=(x-1)^2-(x-3)^2=(x-1-y+3)(x-1+y-3)=(x-y+2)(x+y-4)

OpenStudy (anonymous):

thanks friend !! Medal for you...☺